Given a normal distribution with μ = 180 and σ = 20, find


a. The area above 230


b. The area below 215


c. The area between 185 and 225

1
Expert's answer
2022-04-06T15:53:28-0400

We have a normal distribution, "\mu=180, \sigma=20."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}."


"\text{a. } z_1=\cfrac{230-180}{20}=2.50;\\
P(X>230)=P(Z>2. 50)=\\
=1-P(Z<2.50)="

"=1-0.9938=0.0062" ​(from z-table).


"\text{b. } z_2=\cfrac{215-180}{20}=1.75;\\
P(X<215)=P(Z<1.75)=0.9599"

 ​(from z-table).


"\text{c. } z_3=\cfrac{185-180}{20}=0.25;\\
z_4=\cfrac{225-180}{20}=2.25;\\
P(185<X<225)=P(0.25<Z<2.25)=\\
P(Z<2.25)-P(Z<0.25)=\\
=0.9878-0.5987=0.3891."

 ​(from z-table).